2. Prove that if n > 1, then 1(1!) + 2(2!) + ... + n(n!) = (n + 1)! - 1.
Answer #1

Proof by induction. nel (!) = (1+1)! -1 = 21-1 - 2-1 = 1: Let the result true for n. ire induction hypothesis is 1(!!) + 2 (2